Find the acceleration of our galaxy due to the nearest comparably sized galaxy.The approximate mass of each galaxy is
8 ×10^(11) solar masses and they are separated by 2 million light-years.
Each galaxy has a diameter of 100,000 light years.
(Assume 1 light year =10^(16)m Gravitational constant G =10^(-10)((Nm^(2))/(kg^(2))) and mass of sun =2×10^(30)Kg )
Ans 4×10^(-13)
1
Expert's answer
2020-09-21T08:29:54-0400

According to Newton's second law,


"a=F\/m."

What is the force?


"F=G\frac{m\cdot m}{r^2},"

which gives us


"a=\frac{Gm}{r^2}."

Substitute the following values:

"G=10^{-10}\text{ N}\cdot\text{m}^2/\text{kg}^2,\\
m=8×10^{11}\cdot2×10^{30}=16\cdot10^{41}\text{ kg},\\
r=2×10^6\cdot10^{16}=2\cdot10^{22}\text{ m}."



"a=\frac{10^{-10}\cdot16\cdot10^{41}}{(2\cdot10^{22})^2}=4\cdot10^{-13}\text{ m/s}^2."